# Multiple Queries with Multiple Keys: A Precise Prompt Matching Paradigm for Prompt-based Continual Learning

### v2

@@
Abstract
Continual learning requires machine learning models to continuously acquire new knowledge in dynamic environments while avoiding the forgetting of previous knowledge. Prompt-based continual learning methods effectively address the issue of catastrophic forgetting through prompt expansion and selection. However, existing approaches often suffer from low accuracy in prompt selection, which can result in the model receiving biased knowledge and making biased predictions. To address this issue, we propose the Multiple Queries with Multiple Keys (MQMK) prompt matching paradigm for precise prompt selection. The goal of MQMK is to select the prompts whose training data distribution most closely matches that of the test sample. Specifically, Multiple Queries enable precise breadth search by introducing task-specific knowledge, while Multiple Keys perform deep search by representing the feature distribution of training samples at a fine-grained level. Experiments show that MQMK enhances the prompt matching rate by over 30% in challenging scenarios and achieves state-of-the-art performance on three widely adopted continual learning benchmarks. Once this paper is accepted, we will release the code.
@@ Abstract
1 Introduction
Neural networks have become one of the most important models in the field of machine learning (Rumelhart et al., 1986; LeCun et al., 1998; Graves & Graves, 2012). In real-world scenarios, data may not always be fully accessible from the start, and the dynamic nature of the world continuously introduces new data. This requires neural networks to possess the ability for continual learning (Belouadah et al., 2021; Zhou et al., 2024; De Lange et al., 2021), in order to maintain knowledge updates and adapt to new environments. Additionally, humans do not forget old knowledge when learning new knowledge, making continual learning a key technology for building artificial intelligence that mirrors human learning.
@@ 1 Introduction
Figure 1: (a) Prompts with discrepancies between the training dataset and test samples may provide biased knowledge, leading to biased outputs. (b) Prior prompt selector uses a single query for all tasks and task-level keys. Our prompt selector uses task-level breadth queries and class-level depth keys. In our work, the prompts are involved in the prompt selection process, whereas prior work does not include these.
@@
In the context of continual learning, the data across different learning tasks is often non-independent and identically distributed (non-i.i.d.). Non-i.i.d. data requires the model to continuously adjust its parameters based on new data in order to establish decision boundaries that are more suited to the current data. However, this adaptation can lead to the forgetting of previously learned knowledge. Some studies suggest that unless certain measures are taken, the neural network’s ability to recognize previously encountered classes inevitably deteriorates during the learning process for new tasks, a phenomenon known as catastrophic forgetting (Nguyen et al., 2019; McCloskey & Cohen, 1989; McClelland et al., 1995).
@@
Early continual learning methods design sustainable learning from aspects such as parameters (Kirkpatrick et al., 2017; Li & Hoiem, 2017), data (Rebuffi et al., 2017; Shin et al., 2017), gradients (Lopez-Paz & Ranzato, 2017; Mirzadeh et al., 2020), features (Madaan et al., 2021; Pham et al., 2021), and architecture (Yan et al., 2021; Wang et al., 2022a), training a neural network from scratch, which has yielded promising results. Recent continual learning algorithms (Zhou et al., 2024; Wang et al., 2022c; McDonnell et al., 2024) have started to focus on how to use pre-trained models for continual learning, aiming to leverage the strong generalization ability of pre-trained models to adapt to dynamic downstream tasks.
@@
The prompt-based methods (Wang et al., 2022b, c; Smith et al., 2023; Jung et al., 2023; Gao et al., 2024; Kurniawan et al., 2024), as emerging pre-trained model based approaches, counter the forgetting problem and make reasonable use of the generalization knowledge of pre-trained models in a clever manner. These methods utilize a frozen backbone network, such as Vision Transformer (ViT) (Dosovitskiy, 2020), pre-trained on a large-scale dataset like ImageNet (Russakovsky et al., 2015), and adapt the model to downstream continual learning tasks through Visual Prompt Tuning (VPT) (Jia et al., 2022). A major advantage of these approaches is that task-specific knowledge can be stored in the form of prompts, and as tasks increase, the prompts are continuously expanded, thereby forming a prompt pool that encompasses all the knowledge learned by the model. Existing methods (Wang et al., 2022c, b) employ a query-key matching mechanism for prompt selection. Specifically, this approach uses a pre-trained ViT without prompts as the query function and employs task-specific, learnable parameters as the key for the prompt retrieval. However, a key issue is that the matching mechanism is not sufficiently precise. A prompt trained on a inconsistent data distribution with the test samples may not be effective in assisting inference, as shown in Figure 1(a). From the results, improving the matching rate can effectively boost model performance (see Section 3.3).
@@
To improve the matching rate and the consistency between the prompt and the test samples, we propose the Multiple Queries with Multiple Keys (MQMK) matching paradigm. Unlike the current Single Query-Single Key (SQSK) paradigm, where a single query matches a single key (Wang et al., 2022c, b; Smith et al., 2023), our method extends the queries across all tasks and forms a query pool, with each query corresponding to a task, as shown in Figure 1(b). This query pool, which includes all task-specific prompts and contains task-related knowledge, is used for task-level breadth search. Additionally, we extend the original task-level keys to Multiple Keys in each task to better represent the feature distribution of the training samples for deep search. Our experiments demonstrate that the performance is optimal when Multiple Keys are extended to the class level. Through horizontal expansion of the query-key matching mechanism, the MQMK paradigm significantly improves the matching rate, achieving 32.82% improvement in certain scenario, and delivers State-Of-The-Art (SOTA) performance across three challenging and well-recognized continual learning datasets. Our method and experiments provide strong evidence that, in addition to improving prompt quality, directly enhancing the matching mechanism can effectively improve performance.
@@
    For the first time, we discuss the importance of query-key matching rate in improving model performance in prompt-based methods, defining current approaches as the Single Query-Single Key (SQSK) paradigm and introducing the Multiple Queries with Multiple Keys (MQMK) paradigm, which achieves an improvement of 32.82% in matching rate under certain scenario and reaches SOTA performance.
  2. 2.
@@
    We identify that the SQSK paradigm lacks prompt information in the query, specifically task-specific knowledge, which leads to low query precision. Additionally, in SQSK, the prompt itself is not involved in the process of prompt selection. To address these, we extended the Single Query to Multiple Queries for task-level breadth search.
  3. 3.
    After introducing Multiple Queries, the feature gap between classes within a task becomes further pronounced. A single key can no longer represent the feature distribution of the training samples used for prompt optimization. Therefore, we extend the Single Key to Multiple Keys for deep search and find that the performance is optimal when the key granularity is at the class level.
2 Related Work
@@
Continual Learning. Early continual learning methods can generally be classified into five categories: regularization-based methods (Kirkpatrick et al., 2017; Li & Hoiem, 2017), replay-based methods (Rebuffi et al., 2017; Shin et al., 2017), optimization-based methods (Lopez-Paz & Ranzato, 2017; Mirzadeh et al., 2020), representation-based methods (Madaan et al., 2021; Pham et al., 2021), and architecture-based methods (Yan et al., 2021; Wang et al., 2022a). Regularization-based methods address the forgetting problem by adding explicit regularization terms to balance the old and new tasks. Replay-based methods design strategies to retain important old samples in order to preserve the model’s previous knowledge. Optimization-based methods aim to make the gradients of new and old tasks as independent as possible, avoiding interference in order to prevent forgetting. Representation-based methods enhance the compatibility with new knowledge through meta-learning (Finn et al., 2017; Ravi & Larochelle, 2017) and self-supervised learning (Chen et al., 2020). Architecture-based methods adapt to dynamic task objectives through dynamic networks.
@@
Prompt-based Continual Learning. Prompt-based methods are a type of continuous learning method based on pre-trained models. They leverage task-specific knowledge from the prompt and the generalized knowledge from the pre-trained model, resulting in superior performance. L2P (Wang et al., 2022c) and DualPrompt (Wang et al., 2022b) introduce VPT into continual learning, where prompt vectors can be progressively expanded as tasks increase, thereby avoiding catastrophic forgetting caused by adjusting the same set of parameters for different tasks. L2P proposes a query-key matching approach for instance-wise prompt retrieval, while DualPrompt introduces general prompts to learn knowledge across tasks and provides general knowledge. CODA-P (Smith et al., 2023) introduces a decomposed attention-based prompt querying method, where prompts are combined with varying weights. ESN (Wang et al., 2023) addresses stage interference and performance imbalance by using stage-isolated classifiers, energy normalization, and voting-based inference. EvoPrompt (Kurniawan et al., 2024) addresses prompt selection mismatches and adaptive prompting challenges by using a dynamic, evolving prompt memory system that integrates reference and working prompts through optimal transport and bias adjustment. CPrompt (Gao et al., 2024) employs a random prompt selection training approach to address the issue of prompt inconsistency between training and testing phases.
@@
Figure 2: The average accuracy for four scenarios: when SQSK selects False Prompt and True Prompt, when samples initially with False Prompt are manually replaced with True Prompt, and when all samples use True Prompt.
3 Preliminaries
3.1 Problem Setting
@@
Figure 3: Overall pipeline of MQMK. Training: Directly selects the true prompt based on the label information and inputs it into the backbone to obtain the corresponding query for classification. This process updates the prompt, the key associated with the label and the classifier. Inference: Feeds each prompt into the backbone to generate the corresponding query and selects the most consistent query for classification by matching the task-level queries with the class-level keys.
3.2 Model Tuning and Prompt Pool
@@
Following previous prompt-based continual learning works (Wang et al., 2022c, b; Smith et al., 2023; Gao et al., 2024; Kurniawan et al., 2024), we use ViT as the backbone $f=f_{r}\circ f_{e}$ and a single linear layer $W$ as the classification head, where $f_{e}$ is the embedding layer and $f_{r}$ is the Multiple Self-Attention (MSA) layers. Given an image input $\mathbf{x}$, we first divide it into patches $\mathbf{x}_{p}\in\mathbb{R}^{L\times(S^{2}\times C)}$, where $L$ is the token length, $S$ is patch size and $C$ is the number of channels. And then pass them through $f_{e}$ to obtain corresponding embeddings $\mathbf{x}_{e}\in\mathbb{R}^{L\times D}$, where $D$ is embedding (token) dimension (768). Next, the embeddings $\mathbf{x}_{e}$ and the prompt $P$ are fed into $f_{r}$ for feature extraction to obtain the output $f_{r}(P;\mathbf{x}_{e})\in\mathbb{R}^{L\times D}$. The first token of the output, $f_{r}(P;\mathbf{x}_{e})[0]\in\mathbb{R}^{D}$, which serves as the [class] token, is passed through the classification head $W\in\mathbb{R}^{D\times|\mathcal{Y}^{1}\cup\mathcal{Y}^{2}\cup\cdots\cup\mathcal{Y}^{T}|}$ to obtain the prediction, where $|\mathcal{Y}^{1}\cup\mathcal{Y}^{2}\cup\cdots\cup\mathcal{Y}^{T}|$ represents the total number of classes across all tasks.
@@
To address catastrophic forgetting, L2P and DualPrompt design a paradigm where prompts expand with the tasks. Specifically, as tasks expand, the expert prompts (e-prompts) specific to each task and the general prompts (g-prompts) (Wang et al., 2022b) shared across tasks together form the prompt pool ${\mathbf{P}=\{P_{g},P_{1},P_{2},\cdots,P_{M}\}}$, where $P_{g}\in\mathbb{R}^{L_{g}\times D}$ is the g-prompt, $P_{t}\in\mathbb{R}^{L_{e}\times D}$ is the e-prompt of $t$-th task, $M$ is the pool size, $L_{g}$ is the length of g-prompt and $L_{e}$ is the length of e-prompt. To retrieve prompts, they set a learnable parameter $\mathbf{k}_{t}\in\mathbb{R}^{D}$ for the $t$-th e-prompt and form a prompt pool consisting of key-value pairs, $\mathbf{P}=\{P_{g},(\mathbf{k}_{1},P_{1}),(\mathbf{k}_{2},P_{2}),\cdots,(\mathbf{k}_{M},P_{M})\}$. The features extracted by the ViT without prompts, $f_{r}(\mathbf{x}_{e})[0]\in\mathbb{R}^{D}$, are used as the query $Q$, and the prompt is selected by:
$I=\underset{t\subseteq[1,M]}{\operatorname{argmax}}\cos\langle Q,\mathbf{k}_{t}\rangle,$ (1)
@@
3.3 Prompt Matching Rate
We define the true prompt as a prompt that was trained on the task to which the sample belongs, i.e., $I=t$ for a sample $(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{t}$, which means the sample’s distribution aligns with the distribution of the prompt’s training samples. Our experiments show that the matching rate under the SQSK paradigm is only 45.15% in 10-task Split ImageNet-R (Hendrycks et al., 2021). As shown in Figure 2, if the samples initially with False Prompt are manually replaced with True Prompt, the performance can be improved by 26.62%. Interestingly, the accuracy of the samples initially with False Prompt, after being replaced with True Prompt, is even higher than that of the samples initially with the True Prompt. This suggests that samples selected initially with the False Prompt not only require more task-specific knowledge, but also achieve a higher recognition rate once the true knowledge is incorporated.
4 Method
Our method uses the same model and tuning approach as SQSK, with the innovation of our method lying in the design of keys and queries, as well as their matching mechanism. The overall pipeline of our framework is illustrated in Figure 3.
4.1 The Design Goal
We aim to ensure that the test sample conforms to the distribution of the training samples used for prompt optimization, as this will enable the prompt to effectively handle the test sample. To achieve this goal, we need to address the following two questions:
  1. 1.

### v3

@@
Abstract.
Continual learning requires machine learning models to continuously acquire new knowledge in dynamic environments while avoiding the forgetting of previous knowledge. Prompt-based continual learning methods effectively address the issue of catastrophic forgetting through prompt expansion and selection. However, existing approaches often suffer from low accuracy in prompt selection, which can result in the model receiving biased knowledge and making biased predictions. To address this issue, we propose the Multiple Queries with Multiple Keys (MQMK) prompt matching paradigm for precise prompt selection. The goal of MQMK is to select the prompts whose training data distribution most closely matches that of the test sample. Specifically, Multiple Queries enable precise breadth search by introducing task-specific knowledge, while Multiple Keys perform deep search by representing the feature distribution of training samples at a fine-grained level. Each query is designed to perform local matching with a designated task to reduce interference across queries. Experiments show that MQMK enhances the prompt matching rate by over 30% in challenging scenarios and achieves state-of-the-art performance on three widely adopted continual learning benchmarks. The code is available at https://github.com/DunweiTu/MQMK.
@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3755195ccs: Computing methodologies Online learning settingsccs: Computing methodologies Supervised learning by classificationccs: Computing methodologies Transfer learningccs: Computing methodologies Lifelong machine learning
@@
Figure 1. (a) Prompts inconsistent with test samples may introduce bias. (b) Prior prompt selector uses a single query and task-level keys; our method adopts task-level breadth queries and class-level depth keys, involving prompts in the selection process.
1. Introduction
Neural networks have become one of the most important models in the field of machine learning (Rumelhart et al., 1986; LeCun et al., 1998; Graves and Graves, 2012). In real-world scenarios, data may not always be fully accessible from the start, and the dynamic nature of the world continuously introduces new data. This requires neural networks to possess the ability for continual learning (Belouadah et al., 2021; Zhou et al., 2024; De Lange et al., 2021) to maintain knowledge updates and adapt to new environments. Additionally, humans do not forget old knowledge when learning new knowledge, making continual learning a key technology for building artificial intelligence that mirrors human learning.
@@ 1. Introduction
In the context of continual learning, the data across different tasks is often non-independent and identically distributed (non-i.i.d.). Non-i.i.d. data requires the model to continuously adjust its parameters based on new data to establish decision boundaries more suited to the current data. However, this adaptation can lead to forgetting of previously learned knowledge. Some studies suggest that unless certain measures are taken, the neural network’s ability to recognize previously encountered classes inevitably deteriorates during learning of new tasks, a phenomenon known as catastrophic forgetting (Nguyen et al., 2019; McCloskey and Cohen, 1989; McClelland et al., 1995).
@@
Early continual learning methods design sustainable learning from aspects such as parameters (Kirkpatrick et al., 2017; Li and Hoiem, 2017), data (Rebuffi et al., 2017; Shin et al., 2017), gradients (Lopez-Paz and Ranzato, 2017; Mirzadeh et al., 2020), features (Madaan et al., 2021; Pham et al., 2021), and architectures (Yan et al., 2021; Wang et al., 2022c), training a neural network from scratch, which has yielded promising results. Recent continual learning algorithms (Zhou et al., 2024; Wang et al., 2022b; McDonnell et al., 2024; Tu et al., 2025; Yi, 2024) have started to focus on using pre-trained models to leverage their strong generalization ability for adapting to dynamic downstream tasks.
@@
Prompt-based methods (Wang et al., 2022a, b; Smith et al., 2023; Jung et al., 2023; Gao et al., 2024; Kurniawan et al., 2024; Yang et al., 2024; Feng et al., 2024), as emerging pre-trained model–based approaches, counter forgetting by cleverly reusing the generalization knowledge of pre-trained models. These methods utilize a frozen backbone network, such as Vision Transformer (ViT) (Dosovitskiy, 2020) pre-trained on large-scale datasets like ImageNet (Russakovsky et al., 2015), and adapt to continual learning tasks through Visual Prompt Tuning (VPT) (Jia et al., 2022) or Prefix Tuning (Li and Liang, 2021). A major advantage is that task-specific knowledge is stored in prompts, and as tasks increase, prompts are expanded to form a prompt pool covering all learned knowledge.
Existing methods (Wang et al., 2022b, a) use a query-key matching mechanism for prompt selection: a pre-trained ViT without prompts as the query function, and task-specific learnable parameters as the key. However, the matching is often imprecise. A prompt trained on a distribution inconsistent with test samples may fail to assist inference (fig. 1(a)). Results show that improving the matching rate effectively boosts model performance (see section 1).
@@
To improve the matching rate and consistency between prompt and test samples, we propose the Multiple Queries with Multiple Keys (MQMK) paradigm. Unlike the current Single Query–Single Key (SQSK) paradigm (Wang et al., 2022b, a; Smith et al., 2023), our method builds a query pool—each query corresponding to a task—for task-level breadth search (fig. 1(b)). Additionally, we extend task-level keys to Multiple Keys within each task for deep search. Experiments show performance is best when keys are at class level. Due to knowledge discrepancies among queries, we introduce a local matching mechanism to mitigate mutual interference. To address query overhead under resource constraints, we propose MQMK-Efficient Inference (MQMK-EI), which uses a single enhanced query to match multiple keys. By horizontally expanding the query-key matching mechanism, MQMK achieves up to 32.82% improvement in matching rate and delivers SOTA performance across three recognized continual learning datasets. Our results show that, beyond improving prompt quality, directly enhancing the matching mechanism can effectively improve performance.
@@
    We highlight the importance of query-key matching rate in prompt-based methods, define current approaches as SQSK, and introduce the MQMK paradigm, which achieves 32.82% improvement and SOTA performance.
  2. (2)
@@
    We identify that SQSK lacks task-specific knowledge in the query, leading to low precision, and that the prompt itself is not involved in prompt selection. To address this, we extend Single Query to Multiple Queries for task-level breadth search.
  3. (3)
    After introducing Multiple Queries, class-level feature gaps become more pronounced. A single key cannot fully represent training distributions. Therefore, we extend Single Key to Multiple Keys for deep search and find class-level keys perform best. Finally, we propose a local matching mechanism to further enhance matching precision.
2. Related Work
@@
Continual Learning. Early continual learning methods can generally be classified into five categories: regularization-based methods (Kirkpatrick et al., 2017; Li and Hoiem, 2017), replay-based methods (Rebuffi et al., 2017; Shin et al., 2017), optimization-based methods (Lopez-Paz and Ranzato, 2017; Mirzadeh et al., 2020), representation-based methods (Madaan et al., 2021; Pham et al., 2021), and architecture-based methods (Yan et al., 2021; Wang et al., 2022c). Regularization-based methods address the forgetting problem by adding explicit regularization terms to balance the old and new tasks. Replay-based methods design strategies to retain important old samples in order to preserve the model’s previous knowledge. Optimization-based methods aim to make the gradients of new and old tasks as independent as possible, avoiding interference in order to prevent forgetting. Representation-based methods enhance the compatibility with new knowledge through meta-learning (Finn et al., 2017; Ravi and Larochelle, 2017) and self-supervised learning (Chen et al., 2020). Architecture-based methods adapt to dynamic task objectives through dynamic networks.
@@
Prompt-based Continual Learning. Prompt-based methods are a type of continuous learning method based on pre-trained models. They leverage task-specific knowledge from the prompt and the generalized knowledge from the pre-trained model, resulting in superior performance. L2P (Wang et al., 2022b) and DualPrompt (Wang et al., 2022a) introduce Visual Prompt Tuning (VPT) into continual learning, where prompt vectors help mitigate catastrophic forgetting by providing task-specific conditioning while keeping the backbone frozen. CODA-P (Smith et al., 2023) introduces a decomposed attention-based prompt querying method, where prompts are combined with varying weights. ESN (Wang et al., 2023a) addresses stage interference and performance imbalance by using stage-isolated classifiers, energy normalization, and voting-based inference. EvoPrompt (Kurniawan et al., 2024) addresses prompt selection mismatches and adaptive prompting challenges by using a dynamic, evolving prompt memory system that integrates reference and working prompts through optimal transport and bias adjustment. HiDe-Prompt (Wang et al., 2023b) optimizes intra-task prediction, task identity inference, and task-adaptive prediction, using task-specific prompts and contrastive regularization to overcome sub-optimality in self-supervised pretraining. CPrompt (Gao et al., 2024) employs a random prompt selection training approach to address the issue of prompt inconsistency between training and testing phases.
@@
Figure 2. The average accuracy for four scenarios: when SQSK selects False Prompt and True Prompt, when samples initially with False Prompt are manually replaced with True Prompt, and when all samples use True Prompt (Perfect Match).
3. Preliminaries
3.1. Problem Setting
@@
Figure 3. Overall pipeline of MQMK. Training: Select the true prompt based on the label information and input it into hidden states to obtain the corresponding query for classification. This process updates the prompt, the locally matched key and the classifier. Inference: Feed all prompts into hidden states to generate the query pool and select the most consistent query for classification by locally matching the task-level queries with the corresponding class-level keys.
3.2. Model Tuning and Prompt Pool
Following previous prompt-based continual learning works (Wang et al., 2022b, a; Smith et al., 2023; Gao et al., 2024; Kurniawan et al., 2024), we use ViT as the backbone $f=f_{r}\circ f_{e}$ and a single linear layer $W$ as the classification head, where $f_{e}$ is the embedding layer and $f_{r}$ is the Multiple Self-Attention (MSA) layers. The MSA layers can be written as:
(1) $\operatorname{MSA}\left(\mathbf{h}_{Q},\mathbf{h}_{K},\mathbf{h}_{V}\right)=\operatorname{Concat}\left(\mathrm{h}_{1},\ldots,\mathrm{h}_{m}\right)W^{O}$
(2) $\text{where }\mathrm{h}_{i}=\operatorname{Attention}\left(\mathbf{h}_{Q}W_{i}^{Q},\mathbf{h}_{K}W_{i}^{K},\mathbf{h}_{V}W_{i}^{V}\right),$
@@ 3.2. Model Tuning and Prompt Pool
where $\mathbf{h}_{Q}$, $\mathbf{h}_{K}$, and $\mathbf{h}_{V}$ are the input query, key, and value, respectively. $W^{O}$, $W_{i}^{Q}$, $W_{i}^{K}$, and $W_{i}^{V}$ are projection matrices. $m$ is the number of heads, and $i$ is the index of the layer. The function $\operatorname{Attention}(\cdot,\cdot,\cdot)$ denotes the self-attention mechanism. Given an image input $\mathbf{x}$, we first divide it into patches $\mathbf{x}_{p}\in\mathbb{R}^{L\times(S^{2}\times C)}$, where $L$ is the token length, $S$ is patch size and $C$ is the number of channels. And then pass them through $f_{e}$ to obtain corresponding embeddings $\mathbf{x}_{e}\in\mathbb{R}^{L\times D}$, where $D$ is embedding (token) dimension (768). Next, the embeddings $\mathbf{x}_{e}$ and the prompt $P$ are fed into $f_{r}$ for feature extraction to obtain the output $f_{r}(P;\mathbf{x}_{e})\in\mathbb{R}^{L\times D}$. Specifically, $P$ is split into $P^{K},P^{V}\in\mathbb{R}^{L_{P}/2\times D}$, which are prepended to $\mathbf{h}_{K}$ and $\mathbf{h}_{V}$, respectively which can written by $\operatorname{MSA}\left(\mathbf{h}_{Q},[P^{K};\mathbf{h}_{K}],[P^{V};\mathbf{h}_{V}]\right).$ For simplicity, the layer indices of $P$, $\mathbf{h}_{K}$, and $\mathbf{h}_{V}$ are omitted. The first token of the output, $f_{r}(P;\mathbf{x}_{e})[0]\in\mathbb{R}^{D}$, which serves as the [class] token, is passed through the classification head $W\in\mathbb{R}^{D\times|\mathcal{Y}^{1}\cup\mathcal{Y}^{2}\cup\cdots\cup\mathcal{Y}^{T}|}$ to obtain the prediction, where $|\mathcal{Y}^{1}\cup\mathcal{Y}^{2}\cup\cdots\cup\mathcal{Y}^{T}|$ represents the total number of classes across all tasks.
@@
To address catastrophic forgetting, L2P and DualPrompt design a paradigm where a prompt pool is used to adapt to multiple tasks. Specifically, as tasks expand, the expert prompts (e-prompts) specific to each task and the general prompts (g-prompts) (Wang et al., 2022a) shared across tasks together form the prompt pool $\mathbf{P}=\{P_{g},P_{1},P_{2},\cdots,P_{M}\}$}, where $P_{g}\in\mathbb{R}^{L_{g}\times D}$ is the g-prompt, $P_{t}\in\mathbb{R}^{L_{e}\times D}$ is the e-prompt of $t$-th task, $M$ is the pool size, $L_{g}$ is the length of g-prompt and $L_{e}$ is the length of e-prompt. To retrieve prompts, they set a learnable parameter $\mathbf{k}_{t}\in\mathbb{R}^{D}$ for the $t$-th e-prompt and form a prompt pool consisting of key-value pairs, $\mathbf{P}=\{P_{g},(\mathbf{k}_{1},P_{1}),(\mathbf{k}_{2},P_{2}),\cdots,(\mathbf{k}_{M},P_{M})\}$. The features extracted by the ViT without prompts, $f_{r}(\mathbf{x}_{e})[0]\in\mathbb{R}^{D}$, are used as the query $Q$, and the prompt is selected by:
(3) $I=\underset{t\subseteq[1,M]}{\operatorname{argmax}}\cos\langle Q,\mathbf{k}_{t}\rangle,$
@@
3.3. Prompt Matching Rate
We define the true prompt as a prompt that was trained on the task to which the sample belongs, i.e., $I=t$ for a sample $(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{t}$, which means the sample’s distribution aligns with the distribution of the prompt’s training samples. Our experiments show that the matching rate under the SQSK paradigm is only 45.15% in 10-task Split ImageNet-R (Hendrycks et al., 2021). As shown in fig. 2, if the samples initially with False Prompt are manually replaced with True Prompt, the performance can be improved by 26.62% (from 62.06% to 88.68%) with the help of the task ID label. This indicates the significant impact that prompt selection has on performance. Interestingly, the accuracy of the samples initially with False Prompt, after being replaced with True Prompt, is even higher than that of the samples initially with the True Prompt. This phenomenon may be attributed to the fact that the tasks associated with samples selected by the False Prompt are more challenging to identify, whereas category prediction within those tasks is relatively easier. By replacing the False Prompt with the True Prompt, implicit task information is introduced, which leads to improved accuracy—even surpassing that of the samples originally selected by the True Prompt. Perfect matching refers to an idealized setting where all samples are matched with their corresponding True Prompt. This setting can be viewed as the performance upper bound achievable by prompt selection methods. For more detailed settings and explanations of this experiment, please refer to section 1 of supplementary materials.
4. Method
Our method uses the same model and tuning approach as SQSK, with the innovation of our method lying in the design of keys and queries, as well as their learning and matching mechanism. The overall pipeline of our framework is illustrated in fig. 3.
4.1. The Design Goal
We aim to ensure that the test sample follows to the distribution of the training samples used for prompt optimization, as this will enable the prompt to effectively handle the test sample. To achieve this goal, we need to address the following three questions:
  1. (1)